OpenStudy (anonymous):

How do you work with radicals? Like my practice book is asking for the square root of 75. do you just break it down to the lowest simple amounts? like 25 and 3 or do I have to make it lower? and if so, why is there a whole number infront of the square root?

OpenStudy (anonymous):

You are on the right path. A square root means there is a number multiplied by itself that equals the number under the radical. For example \[\sqrt{36} = 6\] because 6*6 = 36. Sometimes they don't simplify that much. With your problem you with still have a \[\sqrt{3}\] but do you know what will be in front of it?

OpenStudy (anonymous):

5? so like a \[5\sqrt{3}\] but I don't really understand why its in the front and not behind a square root sign in the first place. Can I not just write \[\sqrt{5} \sqrt{3}\] But it doesn't look right. So, does the five placed right in front of the sqrt signify it is to be multiplied by the 3?

OpenStudy (anonymous):

Yes and no. \[\sqrt{5}\] is not the same as 5. \[\sqrt{25}=5\] it is in front of the radical because you have already taken the square root of a number to find the 5. Where it is signifies that it is \[5*\sqrt{3}\] does that explain it a little better?

OpenStudy (anonymous):

yes it does. thank you so much.

OpenStudy (anonymous):

Here is a more detailed walk through that I think should help you. http://www.purplemath.com/modules/radicals.htm